8b-4(b-2)=24
what is the value of b?

Answer by atif.muhammad(135)

8b-4(b-2)=24 Let's multiply out the brackets
8b - 4b +8 = 24 Remember -2 x -4 = +8 (a common mistake is that -2 x -4 = -8)
4b +8 = 24
4b = 16
b = 4